Whaqt is the remainder when 2008^2007 is divided by 5
Answer by scott8148(6628)

look at the units digit for powers of 8
8^1=8
8^2=64
8^3=512
8^4=4096
8^5=32768 ___ same units digit as 8^1 ___ the units digit repeats in a 4 number pattern 8-4-2-6

2007 divided by 4 leaves a remainder of 3 ___ so the units digit of 2008^2007 is 2
___ this is the remainder when 2008^2007 is divided by 5